Using the definition of the scalar product, find the angle between the following vectors. (Find the smallest nonnegative angle.) (a) A = 7i − 8j and B = 3i − 5j °

(b) A = −3i + 6j and B = 3i − 4j + 2k  °
Answer #1

a) definition of scalar product :

A .B   = |A| |B| cos@


(7i - 8j) . ( 3i - 5j) = ( sqrt(7^2 + 8^2 ) sqrt(3^2 + 5^2) ) cos@

61 = ( 10.63 ) ( 5.83) cos@

cos@ =0.984


@ = 10.17 deg

b)

A .B   = |A| |B| cos@


(-3i + 6j) . ( 3i - 4j + 2k) = ( sqrt(3^2 + 6^2 ) sqrt(3^2 + 4^2 + 2^2) ) cos@

-33 = (6.71) (5.39) cos@

@ = 155.96 deg